# Some generalizations of Camina Pairs and orders of elements in cosets

### v1

@@ 2. $(2)$
Assume that $U>1$. Then $(N,U)$ satisfies condition $\mathrm{(O)}$. Also, by Claim 1, either $O^{2}(U)\trianglelefteq N$ and $N/O^{2}(U)\text{ is a }2\text{-group},$ or $U$ is solvable. By Lemma 5.3, the former case cannot happen. Hence $U$ is solvable. Observe that $U<H$ since $\langle H^{G}\rangle=G$ so $H\not\leqslant N$. If $H/U$ is of odd order, then it is solvable and thus $H$ is solvable, a contradiction. Hence $H/U$ is of even order. Thus we can find a $2$-element $y\in H\setminus U$. We have $xy\in xH,$ so $xy$ has odd order. However, in the quotient group $G/N$, $N(xy)=Ny$ is a nontrivial $2$-element, which is a contradiction.
Claim $9$. For every $\sigma\in\operatorname{Aut}(G)$, the pair $(G,\sigma(H))$ satisfies condition $\mathrm{(O)}$. Consequently, if $x\in\Delta_{H}(G)$ has odd order, then $x^{G}y^{G}$ contains only odd order elements for every $y\in H.$ Furthermore, if $x\in\Delta_{H}(G)$ has even order, then $x^{G}y^{G}$ contains only even order elements for every $y\in H.$ Recall that $\Delta_{H}(G)=G\setminus\cup_{g\in G}H^{g}$.
@@
Let $x\in\Delta_{H}(G)$ be of odd order. Then $x^{G}\cap H=\emptyset$. Let $y\in H$. Suppose by contradiction that there exists an element $z\in x^{G}y^{G}$ which has even order. Then $z=x^{u}y^{v}$ for some $u,v\in G.$ By the previous claim, the pair $(G,H^{v})$ satisfies condition $\mathrm{(O)}$. Since $x\in\Delta_{H}(G)$, $x^{u}\in G\setminus H^{v}$ has odd order. As $y^{v}\in H^{v}$, condition $\mathrm{(O)}$ implies that $z=x^{u}y^{v}\in x^{u}H^{v}$ has odd order, which is a contradiction. Therefore, $x^{G}y^{G}$ consists of entirely odd order elements for every $x\in\Delta_{H}(G)$ and $y\in H.$ For the remaining claim, the proof is similar.
Claim $10$. Let $x\in\Delta_{H}(G)$ and $h\in H$. Suppose that either $o(x)$ is odd and $o(h)$ is even or $o(x)$ is even and $o(h)$ is odd. Then $x^{G}h^{G}\subseteq\Delta_{H}(G)$.
First, assume that $o(x)$ is odd and $o(h)$ is even. By way of contradiction, suppose that $x^{a}h^{b}\not\in\Delta_{H}(G)$ for some $a,b\in G.$ Then $(x^{a}h^{b})^{g}\in H$ for some $g\in G.$ Thus $z=x^{u}h^{v}\in H$ for some $u,v\in G.$ It follows that $(x^{-1})^{u}z=h^{v}$. Since $(x^{-1})^{u}\in\Delta_{H}(G)$ has odd order and $z\in H$, by Claim $9,$ $h^{v}=(x^{-1})^{u}z$ must have odd order, which contradicts the choice of $h.$ Therefore, $x^{G}h^{G}\subseteq\Delta_{H}(G)$ as required. The case when $o(x)$ is even and $o(h)$ is odd can be argued similarly.
Claim $11$. The final contradiction.
@@
(b) Assume that every derangement in $\Delta_{H}(G)$ has even order and fix $x\in\Delta_{H}(G)$. As $H$ is nonsolvable, $H$ is not a $2$-group and thus $|H|$ is divisible by an odd prime $p$. By Cauchy’s theorem, let $h\in H$ be an element of order $p$. As in the previous case, let $D=x^{G}$ and $C=h^{G}.$ By Claims 9 and 10, $DC\subseteq\Delta_{H}(G)$ and it contains only even order derangements. Hence for each integer $k\geq 1$, the product $DC^{k}$ contains only even order derangements. By [ASH, Theorem 1.1], we have $C^{m}=G$ for some positive integer $m$ and thus $DC^{m}=G\subseteq\Delta_{H}(G)$, which leads to a contradiction as before. This contradiction completes the proof of the claim. ∎
@@
Proof of Corollary 5.
Let $G$ be a counterexample to the corollary with minimal order. Then there exists a $p$-element $x\in G$ such that $xy$ is $p$-regular for all nontrivial $p$-regular elements $y\in G$, but $x\not\in O_{p}(G).$ Let $L=\langle x^{G}\rangle$. Suppose $L<G$. Then by the minimality of $|G|$, $x\in O_{p}(L)$. As $L\unlhd G$, we have $O_{p}(L)\subseteq O_{p}(G)$ and thus $x\in O_{p}(G)$, a contradiction. Hence, we may assume $G=\langle x^{G}\rangle.$
Let $1<N\unlhd G$. Then $xN$ is a $p$-element in $G/N$, and for all nontrivial $p$-regular elements $yN\in G/N$, we can find a nontrivial $p$-regular element $g\in G$ such that $yN=gN$ and thus $xyN=xgN$ is $p$-regular in $G/N$. Therefore, the pair $(G/N,xN)$ satisfies the hypothesis of the corollary, and by the minimality of $|G|$, $xN\in O_{p}(G/N)$. Suppose $O_{p}(G)>1$. Then since $O_{p}(G/O_{p}(G))$ is trivial and by the previous claim $xO_{p}(G)$ lies in $O_{p}(G/O_{p}(G))$, which forces $x\in O_{p}(G)$, a contradiction. So, we may assume $O_{p}(G)=1$.

### v2

@@ 2. $(2)$
Assume that $U>1$. Then $(N,U)$ satisfies condition $\mathrm{(O)}$. Also, by Claim 1, either $O^{2}(U)\trianglelefteq N$ and $N/O^{2}(U)\text{ is a }2\text{-group},$ or $U$ is solvable. By Remark 5.3, the former case cannot happen. Hence $U$ is solvable. Observe that $U<H$ since $\langle H^{G}\rangle=G$ so $H\not\leqslant N$. If $H/U$ is of odd order, then it is solvable and thus $H$ is solvable, a contradiction. Hence $H/U$ is of even order. Thus we can find a $2$-element $y\in H\setminus U$. We have $xy\in xH,$ so $xy$ has odd order. However, in the quotient group $G/N$, $N(xy)=Ny$ is a nontrivial $2$-element, which is a contradiction.
Claim $8$. For every $\sigma\in\operatorname{Aut}(G)$, the pair $(G,\sigma(H))$ satisfies condition $\mathrm{(O)}$. Consequently, if $x\in\Delta_{H}(G)$ has odd order, then $x^{G}y^{G}$ contains only odd order elements for every $y\in H.$ Furthermore, if $x\in\Delta_{H}(G)$ has even order, then $x^{G}y^{G}$ contains only even order elements for every $y\in H.$ Recall that $\Delta_{H}(G)=G\setminus\cup_{g\in G}H^{g}$.
@@
Let $x\in\Delta_{H}(G)$ be of odd order. Then $x^{G}\cap H=\emptyset$. Let $y\in H$. Suppose by contradiction that there exists an element $z\in x^{G}y^{G}$ which has even order. Then $z=x^{u}y^{v}$ for some $u,v\in G.$ By the previous claim, the pair $(G,H^{v})$ satisfies condition $\mathrm{(O)}$. Since $x\in\Delta_{H}(G)$, $x^{u}\in G\setminus H^{v}$ has odd order. As $y^{v}\in H^{v}$, condition $\mathrm{(O)}$ implies that $z=x^{u}y^{v}\in x^{u}H^{v}$ has odd order, which is a contradiction. Therefore, $x^{G}y^{G}$ consists entirely of odd order elements, for every odd order element $x\in\Delta_{H}(G)$ and $y\in H.$ For the remaining claim, the proof is similar.
Claim $9$. Let $x\in\Delta_{H}(G)$ and $h\in H$. Suppose that either $o(x)$ is odd and $o(h)$ is even or $o(x)$ is even and $o(h)$ is odd. Then $x^{G}h^{G}\subseteq\Delta_{H}(G)$.
First, assume that $o(x)$ is odd and $o(h)$ is even. By way of contradiction, suppose that $x^{a}h^{b}\not\in\Delta_{H}(G)$ for some $a,b\in G.$ Then $(x^{a}h^{b})^{g}\in H$ for some $g\in G.$ Thus $z=x^{u}h^{v}\in H$ for some $u,v\in G.$ It follows that $(x^{-1})^{u}z=h^{v}$. Since $(x^{-1})^{u}\in\Delta_{H}(G)$ has odd order and $z\in H$, by Claim $8,$ $h^{v}=(x^{-1})^{u}z$ must have odd order, which contradicts the choice of $h.$ Therefore, $x^{G}h^{G}\subseteq\Delta_{H}(G)$ as required. The case when $o(x)$ is even and $o(h)$ is odd can be argued similarly.
Claim $10$. The final contradiction.
@@
(b) Assume that every derangement in $\Delta_{H}(G)$ has even order and fix $x\in\Delta_{H}(G)$. As $H$ is nonsolvable, $H$ is not a $2$-group and thus $|H|$ is divisible by an odd prime $p$. By Cauchy’s theorem, let $h\in H$ be an element of order $p$. As in the previous case, let $D=x^{G}$ and $C=h^{G}.$ By Claims 8 and 9, $DC\subseteq\Delta_{H}(G)$ and it contains only even order derangements. Hence for each integer $k\geq 1$, the product $DC^{k}$ contains only even order derangements. By [ASH, Theorem 1.1], we have $C^{m}=G$ for some positive integer $m$ and thus $DC^{m}=G\subseteq\Delta_{H}(G)$, which leads to a contradiction as before. This contradiction completes the proof of the claim. ∎
@@
Proof of Theorem 5.
Let $G$ be a counterexample to the theorem with minimal order. Then there exists a $p$-element $x\in G$ such that $xy$ is $p$-regular for all nontrivial $p$-regular elements $y\in G$, but $x\not\in O_{p}(G).$ Let $L=\langle x^{G}\rangle$. Suppose $L<G$. Then by the minimality of $|G|$, $x\in O_{p}(L)$. As $L\unlhd G$, we have $O_{p}(L)\subseteq O_{p}(G)$ and thus $x\in O_{p}(G)$, a contradiction. Hence, we may assume $G=\langle x^{G}\rangle.$
Let $1<N\unlhd G$. Then $xN$ is a $p$-element in $G/N$, and for all nontrivial $p$-regular elements $yN\in G/N$, we can find a nontrivial $p$-regular element $g\in G$ such that $yN=gN$ and thus $xyN=xgN$ is $p$-regular in $G/N$. Therefore, the pair $(G/N,xN)$ satisfies the hypothesis of the theorem, and by the minimality of $|G|$, $xN\in O_{p}(G/N)$. Suppose $O_{p}(G)>1$. Then since $O_{p}(G/O_{p}(G))$ is trivial and by the previous claim $xO_{p}(G)$ lies in $O_{p}(G/O_{p}(G))$, which forces $x\in O_{p}(G)$, a contradiction. So, we may assume $O_{p}(G)=1$.
@@
Acknowledgment
The authors are grateful to the reviewer for the helpful comments and suggestions.